If you currently have $200,000 in your retirement account, make no new contributions, and can earn 8% (annually), how long will it take you to reach your goal of $1,000,000?

Type or paste question here
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,000,000=200,000*(1.08)^n

(1,000,000/200,000)=(1.08)^n

Taking log on both sides;

log 5=n*log (1.08)

n=log 5/log (1.08)

=20.91 years(Approx).